# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ E.3 Proof of Theorem 3
Note that the minimizer of problem (38) is $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$. Therefore, $y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}$ for $i\in\hat{S}$. The Lagrangian of problem (38) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in\hat{S}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (39)
@@
where the $\xi_{i}$’s are the Lagrangian multipliers. Since $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (38), we have
@@
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},i\in\hat{S},\end{array}\right.$ (40)
@@
where $H^{\prime}_{\gamma,\tau}(\cdot)$ denotes the derivative of $H_{\gamma,\tau}(\cdot)$. Recall that $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin{\hat{S}}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in\hat{S}$. Therefore, it follows from the first line of expression (40) that
@@
$\begin{split}\mathbf{0}&=-\frac{1}{n}\sum_{i\notin\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &\in-\frac{1}{n}\sum_{i\notin\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}.\end{split}$ (41)
Similarly, it can be shown from the second line of display (40) that
$0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}.$ (42)
@@
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}+b,\,i\in\hat{S}.\end{split}$ (43)
We can see from expressions (41), (42) and the third line of expression (40) that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ satisfies the KKT conditions of the convex problem (43). Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the minimizer of problem (43). Moreover, it can be easily seen that $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (43) since $\hat{S}\subseteq{}S_{0}$. This implies that
@@ E.4 Proof of Proposition 5
$0\leq Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)=\sum^{T}_{t=1}\left[G^{\gamma}(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})-G(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})\right]\leq\frac{T}{4}\gamma,$ (44)
@@
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},1\leq t\leq T.\end{split}$ (45)
The Lagrangian of problem (45) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in S_{0,t}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (46)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Let $H^{\prime}_{\gamma,\tau}(\cdot)$ denote the derivative of $H_{\gamma,\tau}(\cdot)$ and $V^{\prime}(\cdot)$ denote the derivative of $V(\cdot)$. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (51),
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\enskip{}i\in S_{0,t}.\end{array}\right.$ (47)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (48)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in S_{0,t}}\xi_{t,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (49)
Therefore, it follows from expression (47) that
@@
Similarly, it can be shown for the rest two displays (48) and (49) that

### v2

@@ 7 Disclosure Statement
Note that the minimizer of problem (S40) is $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$. Therefore, $y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}$ for $i\in\hat{S}$. The Lagrangian of problem (S40) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in\hat{S}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (S41)
@@
where the $\xi_{i}$’s are the Lagrangian multipliers. Since $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (S40), we have
@@
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},i\in\hat{S},\end{array}\right.$ (S42)
@@
where $H^{\prime}_{\gamma,\tau}(\cdot)$ denotes the derivative of $H_{\gamma,\tau}(\cdot)$. Recall that $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin{\hat{S}}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in\hat{S}$. Therefore, it follows from the first line of expression (S42) that
@@
$\begin{split}\mathbf{0}&=-\frac{1}{n}\sum_{i\notin\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &\in-\frac{1}{n}\sum_{i\notin\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}.\end{split}$ (S43)
Similarly, it can be shown from the second line of display (S42) that
$0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}.$ (S44)
@@
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}+b,\,i\in\hat{S}.\end{split}$ (S45)
We can see from expressions (S43), (S44) and the third line of expression (S42) that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ satisfies the KKT conditions of the convex problem (S45). Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the minimizer of problem (S45). Moreover, it can be easily seen that $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (S45) since $\hat{S}\subseteq{}S_{0}$. This implies that
@@
$0\leq Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)=\sum^{T}_{t=1}\left[G^{\gamma}(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})-G(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})\right]\leq\frac{T}{4}\gamma,$ (S46)
@@
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},1\leq t\leq T.\end{split}$ (S47)
The Lagrangian of problem (S47) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in S_{0,t}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (S48)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Let $H^{\prime}_{\gamma,\tau}(\cdot)$ denote the derivative of $H_{\gamma,\tau}(\cdot)$ and $V^{\prime}(\cdot)$ denote the derivative of $V(\cdot)$. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (S53),
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\enskip{}i\in S_{0,t}.\end{array}\right.$ (S49)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (S50)
@@
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in S_{0,t}}\xi_{t,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (S51)
Therefore, it follows from expression (S49) that
@@
Similarly, it can be shown for the rest two displays (S50) and (S51) that